OpenStudy (anonymous):

You invest $1,050.00, buying stock in a company at $10.00/share. The shares increase in value by 13% the first year and decrease by 5% the second year. How much is the total value of your stocks in the second year?

OpenStudy (anonymous):

ok lets determine how many stocks we have purchased so we have $1050 and each stock costs 10 so how many stocks did we purchase?

OpenStudy (anonymous):

1050 diveded by 10= 105 stocks?

OpenStudy (anonymous):

Correct ok so he purchased 105 stocks

OpenStudy (anonymous):

Now the stock started off selling at $10 And then after the first year increased by 13% Ok imma try to show you what i was doing earlier Ok so when it increases by 13% the stock equals: 10+10*.13=10(1+.13)=10(1.13) SO after the second year a single stock is worth 10*1.13=?

OpenStudy (anonymous):

11.3

OpenStudy (anonymous):

how do you decrease by 5%?

OpenStudy (anonymous):

Ok good Then the next year it Decreases by 5% so the stock is equal to: 11.30-(11.30*.05)=11.30(1-.05)=11.30*.95=? Notice that when we increase we add and when we decrease we subtract This should make sense intuitively since we are decreasing by a certain percentage that mean we are removing .. subtracting

OpenStudy (anonymous):

im confused on your numbers in that equation. so 11.30-(11.30*.05)= 10735

OpenStudy (anonymous):

10.735

OpenStudy (anonymous):

yes that is correct Also i wanna try to explain it one more time ... so you understand what I did 11.30-(11.30*.05) Now lets factor out the 11.30 11.30(1-.05) Now lets solve within the bracket 11.30(.95) Now solve 11.30(.95)=10.735

OpenStudy (anonymous):

ok now we have 105 stocks and each stock is worth 10.735 SOoooo how much is our total value of stock?

OpenStudy (anonymous):

105*10.735=1,610.25

OpenStudy (anonymous):

yup :) Voila

OpenStudy (anonymous):

thanks again!

OpenStudy (anonymous):

np